Alice has 8 stars and earns a new one each week. Mack has 2 stars and get 2 each week. How many weeks before they have the same number of stars?
1
Expert's answer
2019-08-27T08:40:48-0400

After "x" weeks Alice and Mack will have "8+x" stars and "2+2x" stars respectively. We need such "x" that "8+x=2+2x" .

Solve the equation "8+x=2+2x"; "8+x-2-x=2+2x-2-x" , that is, "x=6" (weeks).

Answer: 6 weeks.